OpenStudy (anonymous):

solve In(1-x)=4 A.-3 b.3 c. e^4-1 D.1-e^4

OpenStudy (tkhunny):

Have you considered transforming the logarithm to the exponential form? \(log_{b}(a) = c \implies b^{c} = a\) You'll be done in a moment.

OpenStudy (anonymous):

this is new to me idk what to do or whats that

OpenStudy (tkhunny):

Well, then there you go. Do exactly what I suggested. Change to the exponential form. \(ln(x) = log_{e}(x)\)

OpenStudy (anonymous):

itll be C then

OpenStudy (anonymous):

@theyatin ?

OpenStudy (tkhunny):

I'll give you one for free. \(ln(1-x)=4 \implies e^{4} = 1-x\). You take it from there.

OpenStudy (anonymous):

so it is C

OpenStudy (tkhunny):

I don't know. I didn't look at the answers. You tell me what it is. Is "C" correct?

OpenStudy (anonymous):

D is ams take exponetial on both isdes 1-x = e^4 1- e^4 = x